# RLJP: Legal Judgment Prediction via First-Order Logic Rule-enhanced with Large Language Models

### v1

@@ 5 Conclusion
In summary, our proposed RLJP framework introduces three key innovations for LJP: (1) A dynamic rule optimization method that formulates judgment rule optimization with CACL as the process of tree-splitting, effectively addressing fixed rules’ adaptability limitations in complex legal cases; (2) RLJP is a logic-semantic co-reasoning architecture combining lightweight semantic prescreening with FOL judgment rules to strengthen judgment reasoning capabilities; (3) Experimental validation demonstrating superior performance over all baseline methods in LJP tasks, particularly in complex, detailed cases.

### v2

@@ 5 Conclusion
In summary, our proposed RLJP framework introduces three key innovations for LJP: (1) A dynamic rule optimization method that formulates judgment rule optimization with CACL as the process of tree-splitting, effectively addressing fixed rules’ adaptability limitations in complex legal cases; (2) RLJP is a logic-semantic co-reasoning architecture combining lightweight semantic prescreening with FOL judgment rules to strengthen judgment reasoning capabilities; (3) Experimental validation demonstrating superior performance over all baseline methods in LJP tasks, particularly in complex detailed cases.